OpenStudy (anonymous):

write the linear equation y+3=(2/5)(x+9) in standard form

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

y=(2x+3)/(5)

OpenStudy (anonymous):

not it

OpenStudy (anonymous):

Ax + By = C -2/5x + y = 3/5

OpenStudy (anonymous):

not one of my choices